OpenStudy (anonymous):

integral of 54/(x^2+8x+97) dx, x=9tan t-4

OpenStudy (anonymous):

complete the square of the denominator and make a trig substitution

OpenStudy (anonymous):

do you know how to complete the square?

OpenStudy (anonymous):

no i not.

OpenStudy (loser66):

How??? a student study integral but do not know how to complete the square??????

OpenStudy (anonymous):

I knew how to at one point, but it has been awhile.

OpenStudy (anonymous):

ok, we have: x^2+8x+97, but (x+4)^2 = x^2 + 8x + 16, so we need to add more 81 units to get the same thing, so: x^2+8x+97 = (x+4)^2 + 81

OpenStudy (anonymous):

okay makes sense.

OpenStudy (anonymous):

now make a trig substitution, like x+4 = tan^2(u)

OpenStudy (anonymous):

sorry, not this

OpenStudy (anonymous):

call (x+4)/9 = tan^2(u)

OpenStudy (anonymous):

okay

OpenStudy (loser66):

support: let (x+4) = 9tan u

OpenStudy (anonymous):

yes that's right

OpenStudy (loser66):

square both sides: (x+4)^2 =81 tan^2 u

OpenStudy (anonymous):

sorry hahaha

OpenStudy (loser66):

@M4thM1nd continue your stuff, please

OpenStudy (anonymous):

now on is pretty straightforward

OpenStudy (anonymous):

do you know how to continue @lanarose21 ?

OpenStudy (anonymous):

u sub?

OpenStudy (anonymous):

after you complete the square, call (x+4)/9 = tan(u)

OpenStudy (anonymous):

okay. I did that!

OpenStudy (anonymous):

remenber that tan^2(u) + 1 = sec^2(u)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so where are you stuck?

OpenStudy (anonymous):

getting the final answer. like simplifying all of it.

OpenStudy (anonymous):

Refer to the Mathematica attachment.

OpenStudy (anonymous):

Thank you!